# Disentangling Memory and Reasoning Ability in Large Language Models

### v1

@@ 4.2 Main Results
Our algorithm consistently achieves superior performance across multiple benchmarks compared to existing approaches. On StrategyQA, LLaMA-3.1-8B fine-tuned by our method achieves 78.0%, outperforming CoT (69.4%) and Planning-token (76.7%), while LLaMA-2-7B fine-tuned with our method achieves 70.6%, compared to LoRA (61.2%) and Planning-token (63.5%). Similarly, in CommonsenseQA, "Ours" with LLaMA-3.1-8B scores 82.3%, surpassing CoT (70.6%) and Planning-token (76.9%), while LLaMA-2-7B with our method achieves 71.1%, outperforming LoRA (65.4%). A significant breakthrough is observed on the TruthfulQA dataset, where LLaMA-3.1-8B enhanced by our method achieves 86.6%, outperforming GPT-4o by 1.2% in both zero-shot and CoT settings, and also surpassing Planning-token (82.5%) and LoRA (79.8%) in fine-tuned settings. Notably, on StrategyQA, Qwen2.5-7B fine-tuned with our method achieves 78.6%, exceeding Planning-token performance (77.4%) and LoRA (69.1%). However, on the TruthfulQA dataset, Qwen2.5-7B struggles with CoT settings, achieving only 56.7%, highlighting the instability of instruction-tuned models like Qwen2.5-7B on certain datasets. These results highlight the robustness and adaptability of our approach in handling inference-intensive tasks across diverse datasets and models. However, adding a CoT inference step can sometimes decrease performance, as also reported by (Sprague et al., 2024), and the unstable training of instruction-tuned models like Qwen2.5-7B further underscores the limitations of existing approaches compared to our method.
@@ B.2 Example
In this study, we leverage both the inference LLM and knowledge LLM based on GPT-4o to generate a dataset. Specifically, we use the StrategyQA dataset as the source for our generation. The StrategyQA dataset is a question-answering dataset designed to evaluate models’ multi-hop reasoning abilities. It includes questions that require strategic thinking and often demand more than one piece of information to answer correctly. An example is shown below. From this example, we can see that this data generation method perfectly decouples knowledge and reasoning, demonstrating that this approach generates high-quality data.

### v2

@@ 3.2 Main Results
Our algorithm consistently achieves superior performance across multiple benchmarks compared to existing approaches. On StrategyQA, LLaMA-3.1-8B fine-tuned by our method achieves 78.0%, outperforming CoT (69.4%) and Planning-token (76.7%), while LLaMA-2-7B fine-tuned with our method achieves 70.6%, compared to LoRA (61.2%) and Planning-token (63.5%). Similarly, in CommonsenseQA, "Ours" with LLaMA-3.1-8B scores 82.3%, surpassing CoT (70.6%) and Planning-token (76.9%), while LLaMA-2-7B with our method achieves 71.1%, outperforming LoRA (65.4%). A significant breakthrough is observed on the TruthfulQA dataset, where LLaMA-3.1-8B enhanced by our method achieves 86.6%, outperforming GPT-4o by 1.8% in zero-shot setting and 1.2% in CoT setting, and also surpassing Planning-token (82.5%) and LoRA (79.8%) in fine-tuned settings. Notably, on StrategyQA, Qwen2.5-7B fine-tuned with our method achieves 78.6%, exceeding Planning-token performance (77.4%) and LoRA (69.1%). On the TruthfulQA dataset, Qwen2.5-7B struggles with CoT settings, achieving only 56.7%, which underscores the instability of instruction-tuned models like Qwen2.5-7B on certain datasets. These results emphasize the robustness and adaptability of our approach in handling inference-intensive tasks across diverse datasets and models. However, adding a CoT inference step can sometimes decrease performance, as also reported by Sprague et al. (2024). The unstable training of instruction-tuned models like Qwen2.5-7B further highlights the limitations of existing methods compared to our approach.
@@ B.2 Example
In this study, we leverage both the inference LLM and knowledge LLM based on GPT-4o to generate a dataset. Specifically, we use the StrategyQA dataset as the source for our generation. The StrategyQA dataset is a question-answering dataset designed to evaluate models’ multi-hop inference abilities. It includes questions that require strategic thinking and often demand more than one piece of information to answer correctly.